### v2

@@ 2.1 Tariffs and Real Per Capita Consumption
where $\mu_{H}$ denotes the ratio of tariff revenue to labor income and is given by $\mu_{H}=\widetilde{T}_{H}/(w_{H}L_{H})$. [^18] Strictly speaking, the second term in (10), $d\ln\mu_{H}=d\mu_{H}/\mu_{H}$, is not defined at zero tariff (zero revenue). However, the discussion in the text also goes through in this case. An increase in tariffs in Home reduces import demand there and raises Home’s export price relative to its import price to maintain balanced trade. This leads to a higher relative wage in Home, which Matsuyama, (2019) refers to as the terms-of-trade effect in terms of production factors. This mechanism ensures that the first term in (10) is positive, $d\ln w_{H}>0$.
@@
As in the tariff-revenue-to-income ratio $\mu_{H}$, we can show that rising tariffs first increase and then decrease the real per capita consumption $C_{H}/L_{H}$, whether preferences are nonhomothetic CES or homothetic CES including Cobb-Douglas as a special case. [^20] This implies the optimal level of tariff that maximizes welfare. We can analytically derive the optimal uniform tariffs across sectors and show that they are identical regardless of preferences. See Supplemental Appendix A for details. However, an important difference between the two types of preferences lies in the magnitude of the effects of tariffs. Tariffs influence real per capita consumption through the channels we discussed. And this real per capita consumption in turn affects into the aggregate price index. Under nonhomothetic CES preferences, if the income elasticity of manufacturing demand is sufficiently low ($\epsilon^{m}=1<\overline{\epsilon}_{H}$), the rise in consumption driven by tariffs disproportionately expands the high-income-elastic service demand. This pushes up the aggregate expenditure and thus the aggregate price index (i.e., the minimum expenditure per unit of real per capita consumption $P_{H}=E_{H}/(C_{H}/L_{H})$; see footnote 19). The higher aggregate price index lowers welfare gains from tariffs. Therefore, if $\epsilon^{m}=1<\overline{\epsilon}_{H}$, the magnitude of the effect of tariffs on real per capita consumption is smaller under nonhomothetic CES than under homothetic CES. [^21] If the opposite $\epsilon^{m}=1>\overline{\epsilon}_{H}$ holds, in response to the rise in consumption due to tariffs, the low-income-elastic agricultural demand increases disproportionately, pushing down the aggregate expenditure and thus the aggregate price index. Therefore, the magnitude of the effect of tariffs is larger under nonhomothetic CES than under homothetic CES.
@@ 2.2 Tariffs and Expenditure Shares
From Proposition 2, we can infer (not formally prove) how the manufacturing tariff affects the manufacturing expenditure shares in the presence of both nonhomothetic CES preferences and sector-specific tariffs. Suppose Home increases the manufacturing tariff from a sufficiently low level and that $\epsilon^{m}=1<\overline{\epsilon}_{H}$ holds, which is true for developed countries such as the U.S. [^22] The average U.S. tariff on manufacturing products ranged from three to five percent between 1990 and 2014 (see “inward” tariffs in Figure A2 in Supplemental Appendix J for the U.S. and other developed countries). Given our calibrated values in Section 4 ($(\epsilon^{a},\epsilon^{m},\epsilon^{s})=(0.05,1,1.2)$ as in Comin et al.,, 2021), the expenditure-weighted average of nonhomotheticity parameters in the U.S., $\overline{\epsilon}_{US}$, consistently exceeded one and increased in most years from 1965 to 2014 (from 1.06 to 1.13). As the relative price of manufacturing rises, Home shifts expenditure from agriculture and services toward manufacturing. However, the tariff is also likely to increase Home’s real per capita consumption, shifting expenditure from agriculture and manufacturing toward services.
@@ 2.3 Tariffs and Value-added Shares
From Proposition 3, we can infer (not formally prove) the effect of tariffs on sectoral value-added shares in the presence of both nonhomothetic CES preferences and sector-specific tariffs. Suppose Home increases tariffs from a sufficiently low level only in manufacturing sector with a lower income elasticity, $\epsilon^{m}=1<\overline{\epsilon}_{H}$, which seems plausible in developed countries today (see footnote 22). Then the effects of the manufacturing tariff on its value-added share mirror those on the manufacturing expenditure share, captured by the expenditure channel (a) in (12): the relative price effect is likely to contribute positively to the manufacturing value-added share, while the income effect is likely to contribute negatively. In addition, more expensive foreign manufacturing varieties will be replaced with domestic varieties, leading to a higher manufacturing value-added share, which is captured by the net export channel in (12). We will quantify the magnitudes of those three channels in later sections.
@@ 2.4 Toward a Full Dynamic Model
Saving Decisions and a Tariff Shock:   In each period $t\in\{1,2\}$, a representative household in country $H$ earns labor and capital income, $w_{H,t}L_{H,t}+r_{H,t}K_{H,t}$, as well as tariff revenue, $\widetilde{T}_{H,t}$, and spends on consumption, $E_{n,t}$, and investment, $P_{H,t}^{K}I_{n,t}$:
@@ 3 Quantitative Model
where $P_{n,t}^{K}$ is the capital good price index which will be defined later, and $I_{n,t}$ is the quantity of investment. The right-hand side represents the national income, taking into account aggregate trade imbalances. There is a global portfolio that collects an exogenous $\phi_{n,t}$ share of total value added, $w_{n,t}L_{n,t}+r_{n,t}K_{n,t}$, plus tariff revenue, $\widetilde{T}_{n,t}$, and redistributes a per capita transfer, $T_{t}^{P}$, to each country to balance the global portfolio’s budget. In this way, we model trade imbalances as transfers, abstracting from cross-border borrowing and lending (Sposi et al.,, 2024).
@@
where $\Phi_{l}$ is the derivative of $I_{n,t}=\Phi(K_{n,t+1},K_{n,t};\delta_{n,t})$ with respect to its $l$-th argument,252525 Specifically, $\Phi_{1}(K_{n,t+1},K_{n,t})=\frac{\partial\Phi(K_{n,t+1},K_{n,t})}{\partial K_{n,t+1}}=\frac{1}{\lambda}\delta_{n,t}^{1-\frac{1}{\lambda}}\left[\frac{K_{n,t+1}}{K_{n,t}}-(1-\delta_{n,t})\right]^{\frac{1}{\lambda}-1},$ $\Phi_{2}(K_{n,t+1},K_{n,t})=\frac{\partial\Phi(K_{n,t+1},K_{n,t})}{\partial K_{n,t}}=\Phi_{1}(K_{n,t+1},K_{n,t})\cdot\left[(\lambda-1)\frac{K_{n,t+1}}{K_{n,t}}-\lambda(1-\delta_{n,t})\right].$ and
@@
The associated transversality condition is also obtained. Both $E_{n,t+1}/E_{n,t}$ and $\bar{\epsilon}_{n,t+1}/\bar{\epsilon}_{n,t}$ are both increasing in $\frac{C_{n,t+1}/L_{n,t+1}}{C_{n,t}/L_{n,t}}$. Since $\psi>1$, the left-hand side is just an increasing function of the ratio in per capita consumption between periods $t+1$ and $t$. Equation (23) tells that this per capita consumption ratio depends on the discount factor $\beta$, (i) the ratio in the intertemporal demand shifters, (ii) the ratio in populations, and (iii) the real return to capital.
@@
where $\kappa_{n,t}^{j,j^{\prime}}$ is the shifter for sector $j$’s demand for sector $j^{\prime}$ composite good, $M_{n,t}^{j,j^{\prime}}(z)$ is the input of sector $j^{\prime}$ composite good for production of variety $z$ of sector $j$ and is produced using the sectoral composite goods, and $\sigma^{j}$ is the elasticity of substitution across sectoral composite goods for production of sector $j$ varieties. In production of varieties in sector $j$, the cost share of sector $j^{\prime}$ inputs within intermediate-input costs is
@@ 4.1 Data
Our primary data source is the World Input-Output Database (WIOD) Release 2016 and the Long-Run WIOD (Woltjer et al.,, 2021; Timmer et al.,, 2015), which allows us to observe the intermediate input uses across different countries and sectors of both origins and destinations. By merging the two datasets, we constructed a database that spans half a century, 1965–2014. Our empirical exercise encompasses 24 countries (see Table 1) and the rest of the world (RoW). They are the listed countries in the Long-Run WIOD, and we classify Hong Kong as part of the RoW. We aggregate the ISIC industries into three sectors (see Table A4 in Supplemental Appendix E). We complement the WIOD data with the Penn World Table (PWT) 10.0 (Feenstra et al.,, 2015) and CEPII Gravity database (Mayer and Zignago,, 2011). Bilateral tariffs on goods are sourced from the World Integrated Trade Solution (WITS). [^26] As the bilateral tariffs are specified at each Harmonized System (HS) product level, we first group the HS products to the ISIC industries using the concordance table provided by the WITS and then compute the simple average for agriculture and manufacturing sectors.
@@ 4.2 Calibration of Structural Parameters and Fundamentals
We begin with our calibration of the parameters in preferences. The discount factor $\beta$ is set at 0.96 following the literature on macroeconomics. We set the inverse of the intertemporal elasticity of substitution $\psi=2$ following Ravikumar et al., (2019). For parameters in the period utility, we choose the elasticities of substitution across sectors $\sigma=0.5$, and the degree of nonhomotheticity in three sectors is $(\epsilon^{a},\epsilon^{m},\epsilon^{s})=(0.05,1,1.2)$ following Comin et al., (2021). $\sigma<1$ implies that sectoral goods are complements, and, therefore, the relative price effect is at work. The values of $\{\epsilon^{j}\}_{j}$ indicate that the income effect also operates: agriculture is a necessity, services are a luxury, and manufacturing can be either (see (4)).
Shares of primary production factors in the total production cost $\gamma_{n,t}^{j}$ are directly observed in the IO table. Cost shares of capital within value added $\alpha_{n,t}^{j}$ are calibrated as one minus labor shares, which are obtained from the PWT. Since the PWT does not provide the sectoral labor share, we apply the common value across sectors for each year and country. We set the elasticity of substitution across intermediate inputs $\sigma^{j}=0.38$ for all $j$ following Atalay, (2017). For the capital goods production, we set the elasticity of substitution $\sigma^{K}=0.29$ following Sposi et al., (2024). Lower-than-one elasticities of substitution indicate that the relative price effect applies to both the production of goods and services and investment goods. Shape parameters of the Fréchet distribution are calibrated based on the estimates of Caliendo and Parro, (2015). We choose $\theta^{a}=8.11$ and $\theta^{m}=4.55$. For the service trade elasticity, we follow Gervais and Jensen, (2019) and set $\theta^{s}=0.75\cdot\theta^{m}$. We set the adjustment cost elasticity in the low of motion for capital $\lambda=0.75$ following Eaton et al., (2016), and the depreciation rate of capital $\delta_{n,t}$ is obtained from the PWT.
@@ 4.3 Solution Algorithm
We solve the equilibrium transition path backward. We first solve the model for the steady state according to Definition 2, assuming the 2014 fundamentals (e.g., productivity, trade costs, exogenous demand shifters, etc.) remain constant forever. We then suppose that the economy will reach the steady state 450 years after 2014. The solution algorithm for the transition path has two loops: the outer loop finds the sequence of investment (saving) rates, $\{\rho_{n,t}\}_{n,t}$ in (46), that satisfies the dynamic optimality condition governed by the Euler equation (23) and the inner loop solves the intratemporal equilibrium for each period (i.e., solving the sectoral prices and factor prices that satisfy the equilibrium conditions listed in Definition 1). More specifically, for the given sequence of $\{\rho_{n,t}\}_{n,t}$ and the initial period capital stock, we first solve the static equilibrium period-by-period sequentially from 1965 to 2464. After we solve the periodic equilibria up to the year 2464, we update $\rho_{n,t}$ backward from 2464 to 1965 according to the Euler equation. More details are in Supplemental Appendix H (see also Ravikumar et al.,, 2019 for the details of the outer loop iteration).
@@ 5 Counterfactuals
Real Per Capita Consumption and Saving:   Panel (A) of Figure 3 shows the percent changes in real per capita consumption $C_{US,t}/L_{US,t}$ from the baseline to this counterfactual equilibrium. U.S. real per capita consumption is 0.72 percent higher than the baseline equilibrium in 2001. In relation to Proposition 1(i), this result confirms that the U.S. in 2001 was in the increasing part of the tariff–consumption schedule, and a 20 percentage point increase in the tariff is small enough to raise the real consumption. The figure also shows that the difference between the baseline and counterfactual is largest in the first year after the shock, and it diminishes over time. This is because the household has an incentive to front-load the consumption.
Consistent with the intuition discussed in Section 2.4 with a two-country, two-period model, households anticipating the permanent lump-sum transfer of tariff revenue are disincentivised to save, leading to a lower-than-baseline saving rate in the counterfactual. [^27] If an increase in tariffs is temporal (e.g., lasting four years during a U.S. presidency), the implication for savings is opposite. Supplemental Appendix L.1 shows that households anticipating no additional tariff revenue in the future will save more during the high tariff period to smooth out consumption over time. Panel (B) shows that the saving rate, $\rho_{US,t}$ in (46), is approximately 0.2 percentage points lower than the baseline equilibrium. A lower saving rate implies a lower capital stock, resulting in lower real consumption in the long run. While the real per capita consumption remains above the baseline level for the first 60 years after the shock, it falls below the baseline level afterwards.
@@
In terms of magnitude, the sectoral value-added shares respond more to the tariff shock than the sectoral expenditure shares. The manufacturing expenditure share under the counterfactual is approximately 0.2 percentage point higher than the baseline, whereas its value-added share is 0.9–1.3 percentage points higher. This discrepancy is due to our quantitative model featuring international trade as well as input-output linkages through intermediate inputs. [^29] In other words, if we abstract from these two features, the sectoral expenditure shares and value-added shares would be exactly the same as in Comin et al., (2021).
Welfare:   We evaluate the overall welfare implications of the tariff shock measured by consumption equivalent from the viewpoint as of 2001 (see the Supplemental Appendix D for the formula). The second column of Table 1 shows that the U.S. welfare rises by 0.36 percent while all other countries are worse off. [^30] Real per capita consumption of the U.S. will fall below the baseline level in six decades after the shock. Yet, our welfare measures discount the future real consumption, resulting in a positive welfare implication. The imposed tariff directly lowers U.S. imports from other countries, thereby reducing their factor prices and welfare due to the terms-of-trade effect. [^31] This point is discussed using our two-country model in Section 2.1 and Supplemental Appendix A. Among others, Canada, India, and Mexico experience the largest welfare losses (1.26, 0.74, and 0.64 percent, respectively), largely due to the U.S. being their primary trading partner and a major importer of their manufacturing goods.
@@
Australia -0.026 -0.024 -0.048 -0.055
Austria -0.126 -0.151 -0.092 -0.109
Belgium -0.095 -0.110 -0.075 -0.088
Brazil -0.177 -0.195 -0.130 -0.146
Canada -1.260 -1.605 -1.078 -1.385
China -0.361 -0.424 -0.220 -0.255
Germany -0.085 -0.094 -0.060 -0.068
Denmark -0.048 -0.058 -0.041 -0.050
Spain -0.096 -0.112 -0.073 -0.086
Finland -0.141 -0.171 -0.101 -0.122
France -0.069 -0.080 -0.051 -0.059
U.K. -0.064 -0.073 -0.044 -0.051
Greece -0.157 -0.182 -0.126 -0.145
India -0.744 -0.791 -0.442 -0.465
Ireland -0.074 -0.088 -0.067 -0.081
Italy -0.103 -0.120 -0.076 -0.088
Japan -0.070 -0.083 -0.046 -0.053
Korea -0.166 -0.197 -0.134 -0.157
Mexico -0.640 -0.715 -0.716 -0.817
Netherland -0.089 -0.104 -0.072 -0.084
Portugal -0.132 -0.156 -0.109 -0.129
Sweeden -0.108 -0.129 -0.078 -0.093
Taiwan -0.212 -0.250 -0.150 -0.176
U.S. 0.357 0.408 -0.121 -0.165
Rest of the World -0.128 -0.144 -0.078 -0.087
@@
Impacts under Different Preferences:   To highlight the role of nonhomotheticity in accounting for the impacts of tariffs on the sectoral allocation, Figure 4 compares the results for manufacturing expenditure shares in panel (A) and value-added shares in panel (B) across three different preferences; nonhomothetic CES, homothetic CES, and Cobb-Douglas. [^32] The expressions are given in (15) and footnote 24.
@@
Under homothetic CES preferences, both the expenditure share and value-added share respond more in magnitude to the tariff shock than under nonhomothetic CES preferences. This is because homothetic CES preferences shut down the negative income effect on manufacturing shares, which counteracts the positive relative price effect. For the manufacturing expenditure shares presented in panel (A), the gap between the dashed and solid lines, 0.03 percentage point, implies the mitigating impacts coming from the income effect. Under Cobb-Douglas preferences, expenditure shares are determined by the parameters, and therefore, tariffs have no impact.
Panel (B) shows that the manufacturing value-added share is 0.06–0.08 percentage point higher under homothetic CES than under nonhomothetic CES preferences, a magnitude similar to that observed for expenditure shares. However, under Cobb-Douglas preferences where the expenditure-share channel is shut off (i.e., term (a) in (12) vanishes), the model replicates the result very close to that obtained under nonhomothetic CES preferences. This suggests that the primary driver of the effect on value-added shares is a change in the sectoral net exports, corresponding to the term (b) in (12).

### v4

@@ 2.1 Tariffs and Real Per Capita Consumption
where $\mu_{H}$ denotes the ratio of tariff revenue to labor income and is given by $\mu_{H}=\widetilde{T}_{H}/(w_{H}L_{H})$. [^21] Strictly speaking, the second term in (10), $d\ln\mu_{H}=d\mu_{H}/\mu_{H}$, is not defined at zero tariff (zero revenue). However, the discussion in the text also goes through in this case. An increase in tariffs in Home reduces import demand there and raises Home’s export price relative to its import price to maintain balanced trade. This leads to a higher relative wage in Home, which Matsuyama2019 refers to as the terms-of-trade effect in terms of production factors. This mechanism ensures that the first term in (10) is positive, $d\ln w_{H}>0$.
@@
As in the tariff-revenue-to-income ratio $\mu_{H}$, we can show that rising tariffs first increase and then decrease the real per capita consumption $C_{H}/L_{H}$, whether preferences are nonhomothetic CES or homothetic CES including Cobb-Douglas as a special case. [^23] This implies the optimal level of tariff that maximizes welfare. We can analytically derive the optimal uniform tariffs across sectors and show that they are identical regardless of preferences. See Supplemental Appendix A for details. However, an important difference between the two types of preferences lies in the magnitude of the effects of tariffs. Tariffs influence real per capita consumption through the channels we discussed. And this real per capita consumption in turn affects into the aggregate price index. Under nonhomothetic CES preferences, if the income elasticity of manufacturing demand is sufficiently low ($\epsilon^{m}=1<\overline{\epsilon}_{H}$), the rise in consumption driven by tariffs disproportionately expands the high-income-elastic service demand. This pushes up the aggregate expenditure and thus the aggregate price index (i.e., the minimum expenditure per unit of real per capita consumption $P_{H}=E_{H}/(C_{H}/L_{H})$; see footnote 22). The higher aggregate price index lowers welfare gains from tariffs. Therefore, if $\epsilon^{m}=1<\overline{\epsilon}_{H}$, the magnitude of the effect of tariffs on real per capita consumption is smaller under nonhomothetic CES than under homothetic CES. [^24] If the opposite $\epsilon^{m}=1>\overline{\epsilon}_{H}$ holds, in response to the rise in consumption due to tariffs, the low-income-elastic agricultural demand increases disproportionately, pushing down the aggregate expenditure and thus the aggregate price index. Therefore, the magnitude of the effect of tariffs is larger under nonhomothetic CES than under homothetic CES.
@@ 2.2 Tariffs and Expenditure Shares
From Proposition 2, we can infer (not formally prove) how the manufacturing tariff affects the manufacturing expenditure shares in the presence of both nonhomothetic CES preferences and sector-specific tariffs. Suppose Home increases the manufacturing tariff from a sufficiently low level and that $\epsilon^{m}=1<\overline{\epsilon}_{H}$ holds, which is true for developed countries such as the U.S. [^25] The average U.S. tariff on manufacturing products ranged from three to five percent between 1990 and 2014 (see “inward” tariffs in Figure A2 in Supplemental Appendix J for the U.S. and other developed countries). Given our calibrated values in Section 4 ($(\epsilon^{a},\epsilon^{m},\epsilon^{s})=(0.05,1,1.2)$ as in CLM), the expenditure-weighted average of nonhomotheticity parameters in the U.S., $\overline{\epsilon}_{US}$, consistently exceeded one and increased in most years from 1965 to 2014 (from 1.06 to 1.13). As the relative price of manufacturing rises, Home shifts expenditure from agriculture and services toward manufacturing. However, the tariff is also likely to increase Home’s real per capita consumption, shifting expenditure from agriculture and manufacturing toward services.
@@ 2.3 Tariffs and Value-added Shares
From Proposition 3, we can infer (not formally prove) the effect of tariffs on sectoral value-added shares in the presence of both nonhomothetic CES preferences and sector-specific tariffs. Suppose Home increases tariffs from a sufficiently low level only in manufacturing sector with a lower income elasticity, $\epsilon^{m}=1<\overline{\epsilon}_{H}$, which seems plausible in developed countries today (see footnote 25). Then the effects of the manufacturing tariff on its value-added share mirror those on the manufacturing expenditure share, captured by the expenditure channel (a) in (12): the relative price effect is likely to contribute positively to the manufacturing value-added share, while the income effect is likely to contribute negatively. In addition, more expensive foreign manufacturing varieties will be replaced with domestic varieties, leading to a higher manufacturing value-added share, which is captured by the net export channel in (12). We will quantify the magnitudes of those three channels in later sections.
@@ 2.4 Toward a Full Dynamic Model
Saving Decisions and a Tariff Shock: In each period $t\in\{1,2\}$, a representative household in country $H$ earns labor and capital income, $w_{H,t}L_{H,t}+r_{H,t}K_{H,t}$, as well as tariff revenue, $\widetilde{T}_{H,t}$, and spends on consumption, $E_{H,t}$, and investment, $P_{H,t}^{K}I_{n,t}$:
@@ 3 Quantitative Model
where $P_{n,t}^{K}$ is the capital good price index which will be defined later, and $I_{n,t}$ is the quantity of investment. The right-hand side represents the national income, taking into account aggregate trade imbalances. There is a global portfolio that collects an exogenous $\phi_{n,t}$ share of total value added, $w_{n,t}L_{n,t}+r_{n,t}K_{n,t}$, plus tariff revenue, $\widetilde{T}_{n,t}$, and redistributes a per capita transfer, $T_{t}^{P}$, to each country to balance the global portfolio’s budget. In this way, we model trade imbalances as transfers, abstracting from cross-border borrowing and lending (Sposietal2024).
@@
where $\Phi_{l}$ is the derivative of $I_{n,t}=\Phi(K_{n,t+1},K_{n,t};\delta_{n,t})$ with respect to its $l$-th argument, [^28] Specifically, $\Phi_{1}(K_{n,t+1},K_{n,t})=\frac{\partial\Phi(K_{n,t+1},K_{n,t})}{\partial K_{n,t+1}}=\frac{1}{\lambda}\delta_{n,t}^{1-\frac{1}{\lambda}}\left[\frac{K_{n,t+1}}{K_{n,t}}-(1-\delta_{n,t})\right]^{\frac{1}{\lambda}-1},$ $\Phi_{2}(K_{n,t+1},K_{n,t})=\frac{\partial\Phi(K_{n,t+1},K_{n,t})}{\partial K_{n,t}}=\Phi_{1}(K_{n,t+1},K_{n,t})\cdot\left[(\lambda-1)\frac{K_{n,t+1}}{K_{n,t}}-\lambda(1-\delta_{n,t})\right].$ and
@@
The associated transversality condition is also obtained. Both $E_{n,t+1}/E_{n,t}$ and $\bar{\epsilon}_{n,t+1}/\bar{\epsilon}_{n,t}$ are increasing in $\frac{C_{n,t+1}/L_{n,t+1}}{C_{n,t}/L_{n,t}}$. Since $\psi>1$, the left-hand side is just an increasing function of the ratio in per capita consumption between periods $t+1$ and $t$. Equation (23) tells that this per capita consumption ratio depends on the discount factor $\beta$, (i) the ratio in the intertemporal demand shifters, (ii) the ratio in populations, and (iii) the real return to capital.
@@
where $\kappa_{n,t}^{j,j^{\prime}}$ is the shifter for sector $j$’s demand for sector $j^{\prime}$ composite good, $M_{n,t}^{j,j^{\prime}}(z)$ is the input of sector $j^{\prime}$ composite good for production of variety $z$ of sector $j$ and is produced using the sectoral composite goods, and $\sigma^{j}$ is the elasticity of substitution across sectoral composite goods for production of sector $j$ varieties. In sector $j$, the cost share of inputs from sector $j^{\prime}$ within intermediate-input costs is
@@ 4.1 Data
Our primary data source is the World Input-Output Database (WIOD) Release 2016 and the Long-Run WIOD (Woltjer2021; Timmer2015), which allows us to observe the intermediate input uses across different countries and sectors of both origins and destinations. By merging the two datasets, we constructed a database that spans half a century, 1965–2014. Our empirical exercise encompasses 24 countries (see Table 1) and the rest of the world (RoW). They are the listed countries in the Long-Run WIOD, and we classify Hong Kong as part of the RoW. We aggregate the ISIC industries into three sectors (see Table A4 in Supplemental Appendix E). We complement the WIOD data with the Penn World Table (PWT) 10.0 (Feenstra2015) and CEPII Gravity database (Mayer2011). Bilateral tariffs on goods are sourced from the World Integrated Trade Solution (WITS). [^29] As the bilateral tariffs are specified at each Harmonized System (HS) product level, we first group the HS products to the ISIC industries using the concordance table provided by the WITS and then compute the simple average for agriculture and manufacturing sectors.
@@ 4.2 Calibration of Structural Parameters and Fundamentals
We begin with our calibration of the parameters in preferences. The discount factor $\beta$ is set at 0.96 following the literature on macroeconomics. We set the inverse of the intertemporal elasticity of substitution $\psi=2$ following Ravikumar2019. For parameters in the period utility, we choose the elasticities of substitution across sectors $\sigma=0.5$, and the degree of nonhomotheticity in three sectors is $(\epsilon^{a},\epsilon^{m},\epsilon^{s})=(0.05,1,1.2)$ following CLM. $\sigma<1$ implies that sectoral goods are complements, and, therefore, the relative price effect is at work. The values of $\{\epsilon^{j}\}_{j}$ indicate that the income effect also operates: agriculture is a necessity, services are a luxury, and manufacturing can be either (see (4)).
Shares of primary production factors in the total production cost $\gamma_{n,t}^{j}$ are directly observed in the IO table. Cost shares of capital within value added $\alpha_{n,t}^{j}$ are calibrated as one minus labor shares, which are obtained from the PWT. Since the PWT does not provide the sectoral labor share, we apply the common value across sectors for each year and country. We set the elasticity of substitution across intermediate inputs $\sigma^{j}=0.38$ for all $j$ following Atlay2017. For the capital goods production, we set the elasticity of substitution $\sigma^{K}=0.29$ following Sposietal2024. Lower-than-one elasticities of substitution indicate that the relative price effect applies to both the production of goods and services and investment goods. Shape parameters of the Fréchet distribution are calibrated based on the estimates of CaliendoParro2015. We choose $\theta^{a}=8.11$ and $\theta^{m}=4.55$. For the service trade elasticity, we follow Gervais2019 and set $\theta^{s}=0.75\cdot\theta^{m}$. We set the adjustment cost elasticity in the low of motion for capital $\lambda=0.75$ following Eaton2016, and the depreciation rate of capital $\delta_{n,t}$ is obtained from the PWT.
@@ 4.3 Solution Algorithm
We solve the equilibrium transition path backward. We first solve the model for the steady state according to Definition 2, assuming the 2014 fundamentals (e.g., productivity, trade costs, exogenous demand shifters, etc.) remain constant forever. We then suppose that the economy will reach the steady state 450 years after 2014. The solution algorithm for the transition path has two loops: the outer loop finds the sequence of investment (saving) rates, $\{\rho_{n,t}\}_{n,t}$ in (46), that satisfies the dynamic optimality condition governed by the Euler equation (23) and the inner loop solves the intratemporal equilibrium for each period (i.e., solving the sectoral prices and factor prices that satisfy the equilibrium conditions listed in Definition 1). More specifically, for the given sequence of $\{\rho_{n,t}\}_{n,t}$ and the initial period capital stock, we first solve the static equilibrium period-by-period sequentially from 1965 to 2464. After we solve the periodic equilibria up to the year 2464, we update $\rho_{n,t}$ backward from 2464 to 1965 according to the Euler equation. More details are in Supplemental Appendix H (see also Ravikumar2019 for the details of the outer loop iteration).
@@ 5 Counterfactuals
Real Per Capita Consumption and Saving: Panel (A) of Figure 3 shows the percent changes in real per capita consumption $C_{US,t}/L_{US,t}$ from the baseline to this counterfactual equilibrium. U.S. real per capita consumption is 0.72 percent higher than the baseline equilibrium in 2001. In relation to Proposition 1(i), this result confirms that the U.S. in 2001 was in the increasing part of the tariff–consumption schedule, and a 20-percentage-point increase in the tariff is small enough to raise the real consumption. The figure also shows that the difference between the baseline and counterfactual is largest in the first year after the shock, and it diminishes over time. This is because the household has an incentive to front-load the consumption.
Consistent with the intuition discussed in Section 2.4 with a two-country, two-period model, households anticipating the permanent lump-sum transfer of tariff revenue are disincentivised to save, leading to a lower-than-baseline saving rate in the counterfactual. [^30] If an increase in tariffs is temporal (e.g., lasting four years during a U.S. presidency), the implication for savings is the opposite. Supplemental Appendix L.1 shows that households anticipating no additional tariff revenue in the future will save more during the high tariff period to smooth out consumption over time. Panel (B) shows that the saving rate, $\rho_{US,t}$ in (46), is approximately 0.3 percentage points lower than the baseline equilibrium. A lower saving rate implies a lower capital stock, resulting in lower real consumption in the long run. While the real per capita consumption remains above the baseline level for the first 55 years after the shock, it falls below the baseline level afterward.
@@
In terms of magnitude, the sectoral value-added shares respond more to the tariff shock than the sectoral expenditure shares. The manufacturing expenditure share under the counterfactual is approximately 0.2 percentage point higher than the baseline, whereas its value-added share is 0.9–1.3 percentage points higher. This discrepancy is due to our quantitative model featuring international trade as well as input-output linkages through intermediate inputs. [^32] In other words, if we abstract from these two features, the sectoral expenditure shares coincide with the value-added shares as in CLM.
Welfare: We evaluate the overall welfare implications of the tariff shock measured by consumption equivalent from the viewpoint as of 2001 (see the Supplemental Appendix D for the formula). The second column of Table 1 shows that the U.S. welfare rises by 0.41 percent while all other countries are worse off. [^33] Real per capita consumption of the U.S. will fall below the baseline level in six decades after the shock. Yet, our welfare measures discount the future real consumption, resulting in a positive welfare implication. The imposed tariff directly lowers U.S. imports from other countries, thereby reducing their factor prices and welfare due to the terms-of-trade effect. [^34] This point is discussed using our two-country model in Section 2.1 and Supplemental Appendix A. Among others, Canada, India, and Mexico experience the largest welfare losses (1.44, 0.75, and 0.64 percent, respectively), largely due to the U.S. being their primary trading partner and a major importer of their manufacturing goods.
@@
Australia -0.050 -0.054 -0.064 -0.075
Austria -0.139 -0.164 -0.098 -0.116
Belgium -0.123 -0.142 -0.090 -0.105
Brazil -0.187 -0.208 -0.138 -0.155
Canada -1.440 -1.823 -1.208 -1.544
China -0.368 -0.437 -0.226 -0.266
Germany -0.094 -0.105 -0.067 -0.076
Denmark -0.054 -0.065 -0.046 -0.055
Spain -0.106 -0.124 -0.080 -0.094
Finland -0.149 -0.179 -0.106 -0.127
France -0.078 -0.090 -0.057 -0.066
U.K. -0.071 -0.082 -0.050 -0.057
Greece -0.159 -0.184 -0.128 -0.149
India -0.748 -0.800 -0.450 -0.475
Ireland -0.076 -0.089 -0.068 -0.082
Italy -0.107 -0.124 -0.079 -0.091
Japan -0.071 -0.082 -0.046 -0.053
Korea -0.157 -0.184 -0.130 -0.151
Mexico -0.635 -0.706 -0.718 -0.818
Netherland -0.111 -0.129 -0.085 -0.099
Portugal -0.149 -0.174 -0.119 -0.139
Sweeden -0.114 -0.136 -0.082 -0.098
Taiwan -0.211 -0.246 -0.148 -0.172
U.S. 0.408 0.466 -0.138 -0.185
Rest of the World -0.132 -0.149 -0.081 -0.091
@@
Impacts under Different Preferences: To highlight the role of nonhomotheticity in accounting for the impacts of tariffs on the sectoral allocation, Figure 4 compares the results for manufacturing expenditure shares in panel (A) and value-added shares in panel (B) across three different preferences; nonhomothetic CES, homothetic CES, and Cobb-Douglas. [^35] The expressions are given in (15) and footnote 27.
@@
Under homothetic CES preferences, both the expenditure share and value-added share respond more in magnitude to the tariff shock than under nonhomothetic CES preferences. This is because homothetic CES preferences shut down the negative income effect on manufacturing shares, which counteracts the positive relative price effect. For the manufacturing expenditure shares presented in panel (A), the gap between the dashed and solid lines, 0.03–0.04 percentage point, implies the mitigating impacts coming from the income effect. Under Cobb-Douglas preferences, expenditure shares are determined by the parameters, and therefore, tariffs have no impact.
Panel (B) shows that the manufacturing value-added share is 0.05–0.08 percentage point higher under homothetic CES than under nonhomothetic CES preferences, a magnitude similar to that observed for expenditure shares. However, under Cobb-Douglas preferences where the expenditure-share channel is shut off (i.e., term (a) in (12) vanishes), the model replicates the result very close to that obtained under nonhomothetic CES preferences. This suggests that the primary driver of the effect on value-added shares is a change in the sectoral net exports, corresponding to the term (b) in (12).